### v6

# Investigating Memory in RL with POPGym Arcade

Abstract
How should we analyze memory in deep RL? We introduce mathematical tools for fairly analyzing policies under partial observability and revealing how agents use memory to make decisions. To utilize these tools, we present POPGym Arcade [^1] Code available at https://github.com/bolt-research/popgym-arcade, a collection of Atari-inspired, hardware-accelerated, pixel-based environments sharing a single observation and action space. Each environment provides fully and partially observable variants, enabling counterfactual studies on observability. We find that controlled studies are necessary for fair comparisons, and identify a pathology where value functions smear credit over irrelevant history. With this pathology, we demonstrate how out-of-distribution scenarios can contaminate memory, perturbing the policy far into the future, with implications for sim-to-real transfer and offline RL.
1 Introduction
Real-world decision making often requires reasoning under partial observability, where important information is hidden from the agent. This is a challenge for reinforcement learning (RL), which excels in fully observable settings (Silver et al., 2016). To mitigate partial observability, we store and recall sensory information using some form of memory. We often evaluate deep memory models by comparing returns across partially observable tasks. However, deep RL experiments are sensitive to model configurations, observation size, task difficulty, and other confounding factors; especially when combined with memory (Mania et al., 2018; Pleines et al., 2025). This sensitivity makes it difficult to isolate the impact of memory on observability, fairly quantify memory performance, or understand failure modes (Jordan et al., 2024).
To address this gap in understanding, we introduce mathematical tools to isolate and study memory and observability, and to interpret how agents use memory. To utilize these tools, we introduce POPGym Arcade (Fig. 1), an Atari-style hardware-accelerated benchmark where every environment has fully observable and partially observable “twins” with identical spaces and dynamics. By combining twins with our memory analysis tools, we perform counterfactual studies on observability. Our findings motivate a secondary study investigating what memory models learn, where we uncover credit attribution failures.
MDP Environments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
POMDP Environments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Observations from our environment twins. All environments share a unifying observation and action space, enabling counterfactual studies with observability as the sole independent variable.
Contributions
  1. (C1)
    Tools to disentangle, measure, and interpret memory in an RL context.
  2. (C2)
    POPGym Arcade: a benchmark of paired MDP/POMDP twins for controlled studies.
  3. (C3)
    Evidence that memory introduces bias and confounds return-based comparisons.
  4. (C4)
    Discovery of a pathology where memory models incorrectly assign value to irrelevant history.
  5. (C5)
    Demonstration of recurrent state contamination, where C4 causes irrelevant past to corrupt a policy.
2 Preliminaries
Reinforcement Learning
A Markov Decision Process (MDP) (Sutton & Barto, 2018) is a tuple $\mathcal{M}=\langle S,A,P,R,\gamma\rangle$, with the set of states $S$, actions $A$, and transition function $P:S\times A\mapsto\Delta(S)$, where $\Delta(S)$ is a distribution over $S$. MDPs also contain a reward function $R:S\times A\mapsto\Delta(\mathbb{R})$ and discount factor $\gamma\in[0,1)$. At each timestep, we receive a state $s_{t}\in S$ and select an action $a_{t}\in A$ from a policy $a_{t}\sim\pi(\cdot|s_{t})$ that takes us to the next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ emitting a reward $r_{t}\sim R(\cdot|s_{t},a_{t})$. The Markov property states that $s_{t}$ contains sufficient information for an optimal policy. In RL, we learn the policy $\pi:S\mapsto\Delta(A)$ that maximizes the expected return $J(\pi,\mathcal{M})=\mathbb{E}_{\pi,\mathcal{M}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ for $\mathcal{M}$.
POMDPs
In Partially Observable Markov Decision Processes (POMDPs), the agent cannot access environment state, instead perceiving the state through noisy or limited sensors. A POMDP $\mathcal{P}=\langle S,A,P,R,\gamma,\Omega,O\rangle$ extends the MDP with set of observations $\Omega$ and observation function $O:S\mapsto\Delta(\Omega)$. At each timestep, the agent indirectly accesses the state through a non-Markov observation $o_{t}\sim O(s_{t})$. Fortunately, the trajectory $\mathbf{x}_{t}=(o_{0},a_{0},o_{1},a_{1}\dots a_{t-1},o_{t})$, is itself Markov (Sutton & Barto, 2018). The Reward Memory Length (RML) determines how far back in time we must search to learn the optimal policy (Ni et al., 2024). An RML of $k$ expresses that the reward depends on only the latest $k$ elements of the trajectory $r_{t+1}\perp\!\!\!\!\perp\mathbf{x}_{t},a_{t}|\mathbf{x}_{t-k:t},a_{t}$. Our new RL objective depends on a memory model $f$ and POMDP $\mathcal{P}$: $J(f,\pi,\mathcal{P})=\mathbb{E}_{\pi,f,\mathcal{P}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ .
Memory and Recurrences
A memory model $f$ produces a fixed-size latent Markov state $\hat{s}\in\hat{S}$ from the trajectory $\mathbf{x}_{t}$. This confines the complexities associated with partial observability to $f$, allowing the policy to operate on a latent Markov state $\hat{s}$. For conciseness, we introduce the joint observation-action space $X=(\Omega,A);x_{t}=(o_{t},a_{t-1})$. A recurrent memory model $f:X\times H\mapsto\hat{S}\times H$ maintains a summary of the trajectory in the recurrent state $h\in H$ and outputs a latent Markov state $\hat{s}\in\hat{S}$ for the policy
$\hat{s}_{t},h_{t}=f(x_{t},h_{t-1})$ $a_{t}\sim\pi(\cdot\mid\hat{s}_{t}).$ (1)
3 Related Work
POMDP Benchmarks
Progress in RL builds upon benchmarks that provide MDPs and POMDPs (Appendix I). Benchmarks include general-purpose tasks (Bellemare et al., 2013; Morad et al., 2023a) and domain specific tasks like navigation (Chevalier-Boisvert et al., 2018; Beattie et al., 2016; Kempka et al., 2016), psychological experiments (Fortunato et al., 2019), and velocity-masked control tasks (Todorov et al., 2012; Ni et al., 2022), and party games (Pleines et al., 2022). Unfortunately, the cost of training memory models combined with the sample inefficiency of RL is prohibitively expensive, forcing researchers into a tradeoff between simplistic low-dimensional observation spaces and a reduced number of experiments to back their claims. Recent hardware-accelerated benchmarks avoid CPU-GPU copies, resulting in significant training speedups (Hessel et al., 2021; Toledo, 2024), and new high-throughput algorithms like PQN (Gallici et al., 2024). We highlight benchmarks from Matthews et al. (2024a); Pignatelli et al. (2024); Lu et al. (2024); Tao et al. (2025) which offer hardware-accelerated POMDPs.
Controlled Studies of Partial Observability
There is ample literature on the theoretical costs and intractability induced by partial observability (Papadimitriou & Tsitsiklis, 1987; Kaelbling et al., 1998; Vlassis et al., 2012; Liu et al., 2022). However, there is less work that empirically quantifies this performance degradation under deep function approximation. Prior work often compares returns between competing memory models (Parisotto et al., 2020; Morad et al., 2023b; Le et al., 2024). With POMDP returns alone, it is not possible to determine whether greater returns are due to more informative latent Markov states or other factors. Paradoxically, literature shows that in certain cases, using memory in MDPs can improve performance, while using memory in POMDPs can reduce performance (Hausknecht & Stone, 2015; Burda et al., 2022; Morad et al., 2023a). For example, a memory model may increase the parameter count of the model, which can improve policy capabilities (Hansen et al., 2023) without improving the Markov state estimate. On the other hand, optimization of memory-endowed policies is notoriously difficult (Mikhaeil et al., 2022), and a suboptimal memory-free policy may reach better local optima in difficult tasks. To isolate the impact of memory, it is necessary to control for all other variables by comparing performance in a partially observable setting against its fully observable counterpart (Jordan et al., 2024). Ni et al. (2022); Morad et al. (2023a); Rajan et al. (2023) provide capabilities for counterfactual studies, but only Ni et al. (2022); Tao et al. (2025) actually perform them. Few works utilize a single shared observation and action space over all tasks, which is necessary to fully control for both parameter count and task difficulty
Investigating Memory in RL
Comparing returns indirectly measures the information content of the latent Markov state $\hat{s}$ created by the memory model. Few works directly answer how the model constructs the latent Markov state, providing hints but never complete answers. Kapturowski et al. (2019) measure the impact of stale recurrent states. Ni et al. (2024) decouples memory performance from temporal credit assignment, more directly measuring the quality of $\hat{s}$. Finally, Elelimy et al. (2024); Morad et al. (2023b) examine the distribution of recurrent states.
4 Memory Evaluation Tools
The return measures policy performance, but does not disentangle memory capabilities from policy performance, nor does it tell us what the memory model is learning. Below, we describe two tools to isolate and study memory capabilities, and two more tools to understand what memory models learn.
Observability Gap
Consider a base MDP $\mathcal{M}=\langle S,A,P,R,\gamma\rangle$. We can induce partial observability by introducing an observation function $o\sim O(s)$, thus creating a corresponding POMDP twin $\mathcal{P}=\langle\mathcal{M},\Omega,O\rangle$. We are interested in a controlled setting where the observation stream is theoretically sufficient to recover the underlying state, ensuring that the optimal achievable return is identical in both cases.
Our work focuses on the practical difficulty of learning a memory model $f$, to approximate the latent Markov state. To isolate this challenge, we compare the performance of an agent configuration $(\pi,f)$ on $\mathcal{M}$ and $\mathcal{P}$. Under $\mathcal{M}$, $f$ must simply propagate states to the policy. The case for $\mathcal{P}$ is more difficult: the agent must infer the Markov state from the trajectory. Any suboptimality in the learned memory model $f$ will result in a performance gap between the two.
Definition 4.1 (Observability Gap).
The Observability Gap measures the performance loss of a memory-based agent $(f,\pi)$ attributable to inferring the state from observations instead of observing it directly
$\text{Gap}(f,\pi,\mathcal{M},\mathcal{P})=J(f,\pi,\mathcal{M})-J(f,\pi,\mathcal{P})\quad\text{where}\quad\mathcal{M}=\langle S,A,P,R,\gamma\rangle,\quad\mathcal{P}=\langle\mathcal{M},\Omega,O\rangle.$ (2)
Memory Bias
As stated in Section 3, prior work hints that introducing memory can produce confounding factors. To measure the inherent performance cost of introducing memory, we propose the Memory Bias. This is the observed performance difference between a memory-endowed policy and its memory-free counterpart on a fully observable task. Factors that cause Memory Bias could include parameter count, optimization complexity, implicit regularization, representational capacity, and more.
Definition 4.2.
The Memory Bias measures the performance change caused by the memory model that is attributable to the memory model but not attributable to partial observability
$\text{Bias}(f,\pi,\mathcal{M})=J(f,\pi,\mathcal{M})-J(\pi,\mathcal{M}).$ (3)
Pixel Visualizations
Pixel states and observations enable powerful visualization tools. We implement visualization tools to probe which pixels persist in memory via their impact on the policy. More formally, given a trajectory $\textbf{x}_{n}$, we compute latent Markov states $\hat{s}_{0},\hat{s}_{1},\dots,\hat{s}_{n}$ via Eq. 1. Then, we backpropagate through memory and policy, taking the absolute value of the gradient of $Q$ values with respect to an observation $o_{t}$ where $t\leq n$. We provide variants for both $Q$ learning and policy gradient methods
$\sum_{a_{n}\in A}\left\lVert\nabla_{o_{t}}Q(\hat{s}_{n},a_{n})\right\rVert_{2}^{2}=\sum_{a_{n}\in A}\left\lVert\frac{\partial Q(\hat{s}_{n},a_{n})}{\partial\hat{s}_{n}}\frac{\partial\hat{s}_{n}}{\partial o_{t}}\right\rVert_{2}^{2}\quad\int_{A}\left\lVert\nabla_{o_{t}}\pi(a_{n}|\hat{s}_{n})\right\rVert_{2}^{2}\,da_{n}=\int_{A}\left\lVert\frac{\partial\pi(a_{n}|\hat{s}_{n})}{\partial\hat{s}_{n}}\frac{\partial\hat{s}_{n}}{\partial o_{t}}\right\rVert_{2}^{2}\,da_{n}.$
Eq. 4 measures how much each pixel from a prior observation $o_{t}$ propagates through memory and contributes to either the $Q$ values or action distribution at time $n$. We experimented with other norms, but we find the $L_{2}$ norm provides the clearest pixel visualizations.
Recall Density
Eq. 4 produces qualitative saliency maps, describing pixel-level memory recall over a single trajectory. We are also interested in a more quantitative and consistent measure that elucidates the inner workings of memory in expectation over a distribution of trajectories. For these cases, we introduce the recall density, which we derive below. We provide variants for both Q learning and policy gradient. We start with Eq. 4 and reduce over the observation dimensions using the $L_{1}$ norm, such that we receive a scalar time-varying function $z$
$z_{Q}(\mathbf{x}_{n},t)=\sum_{a_{n}\in A}\left\lVert\nabla_{o_{t}}Q(\hat{s}_{n},a_{n})\right\rVert_{1}$ $z_{\pi}(\mathbf{x}_{n},t)=\int_{A}\left\lVert\nabla_{o_{t}}\pi(a_{n}|\hat{s}_{n})\right\rVert_{1}da_{n},$ (5)
where $z_{Q}$ represents the quantity for $Q$ functions and $z_{\pi}$ for policies. Then, we normalize $z$, representing a scale-invariant density over trajectory $\mathbf{x}_{n}$
$\delta_{Q}(\mathbf{x}_{n},t)=\frac{z_{Q}(\mathbf{x}_{n},t)}{\sum_{i=0}^{n}z_{Q}(\mathbf{x}_{n},i)}$ $\delta_{\pi}(\mathbf{x}_{n},t)=\frac{z_{\pi}(\mathbf{x}_{n},t)}{\sum_{i=0}^{n}z_{\pi}(\mathbf{x}_{n},i)}.$ (6)
Finally, we take the sample mean over $m$ trajectories to approximate the expected density under a policy $\pi$ and memory model $f$
$\frac{1}{m}\sum_{i=1}^{m}\delta_{Q}(\mathbf{x}^{(i)}_{n},t)\approx\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x}_{n},t)]$ $\frac{1}{m}\sum_{i=1}^{m}\delta_{\pi}(\mathbf{x}^{(i)}_{n},t)\approx\mathbb{E}_{\pi,f}[\delta_{\pi}(\mathbf{x}_{n},t)].$ (7)
We may compute the resulting expectation over all $0\leq t\leq n$ to generate a plot summarizing memory recall.
Thus far, $n$ must be equal across all sampled trajectories, while in practice trajectories often differ in length. We use generalized time coordinates to resolve this issue. For each trajectory $\mathbf{x}^{(i)}$ of length $n_{i}$, we normalize the timestep $t\in\{0,\dots,n_{i}\}$ to a value $\tau=t/n_{i}\in[0,1]$. We then compute the sample mean of the empirical densities at corresponding normalized time points across $m$ trajectories
$\frac{1}{m}\sum_{i=1}^{m}\delta_{Q}(\mathbf{x}^{(i)}_{n_{i}},\lfloor\tau\cdot n_{i}\rfloor)\approx\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x},\tau)]$ $\frac{1}{m}\sum_{i=1}^{m}\delta_{\pi}(\mathbf{x}^{(i)}_{n_{i}},\lfloor\tau\cdot n_{i}\rfloor)\approx\mathbb{E}_{\pi,f}[\delta_{\pi}(\mathbf{x},\tau)].$ (8)
We may compute the expectation over all $\tau\in[0,1]$ to generate a plot summarizing memory recall. One should ensure that $\mathbf{x}$ is not terminal, or risk evaluating $Q$ or $\pi$ at a terminal state.
Definition 4.3 (Recall Density).
The Recall Density measures the expected relative influence of a past observation at the current timestep. For trajectories generated by policy $\pi$ and memory model $f$, it quantifies how much an observation at normalized time $\tau\in[0,1]$ impacts either:
  1. (4.3.1)
    $\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x},\tau)]$: The $Q$ value at trajectory end, via $Q$ gradient magnitudes
  2. (4.3.2)
    $\mathbb{E}_{\pi,f}[\delta_{\pi}(\mathbf{x},\tau)]$: The action distribution at trajectory end, via policy gradient magnitudes
5 POPGym Arcade
We cannot utilize our memory analysis tools with most prior benchmarks as the Observability Gap requires MDP/POMDP twins and pixel visualizations require both pixel states and observations. To this end, we propose POPGym Arcade (Tables 1, G and H). POPGym Arcade consists of ten base environments each with three difficulty levels. All environments are inspired by existing human-playable card games, board games, and video games, with each environment testing different aspects of memory. Unlike Atari, our tasks utilize (1) highly stochastic transition functions (2) hardware acceleration (3) known RML (4) standardized returns to enable easy comparisons.
Environment Twins
Our twins introduce two notions of state: a low dimensional hidden Markov state $\tilde{s}\in\tilde{S}$ and a pixel Markov state $s\in S$. For example, in MDP MineSweeper, $\tilde{s}$ contains the location of all mines, while $s$ contains pixels representing numbered tiles inferring mine locations. Both states are Markov, but deterministic transitions in $\tilde{S}$ become stochastic in $S$. While $\tilde{S}$ varies across tasks, $S$ is identical for all tasks. By decoupling $\tilde{s}$ from $s$, we can introduce POMDP variants with an observation function $O:\tilde{S}\mapsto\Delta(\Omega)$. Consequently, twins share identical underlying transition dynamics $\tilde{S}\times A\mapsto\tilde{S}$. All tasks share a singular state/observation space $S=\Omega$ and action space $A$. As a result, we can reuse a single model across all tasks and task configurations. We can even change from POMDP to MDP mid-episode, or train a single policy on both MDP and POMDP at the same time. The motivation behind this task structure is to isolate and study memory and observability.
Table 1: A summary of our environments. We provide both small (S) and large (L) state/observation spaces for each task. We describe RML over episode length $n$ or a constant length $k$. Frame stacking or fixed-window attention are sufficient to solve $O(k)$ RML POMDPs but not $O(n)$ POMDPs.
Base Envs Total Envs State/Observation Size Action Space Return Bounds
10 120 S: $128\times 128\times 3$ L: $256\times 256\times 3$ $\uparrow,\downarrow,\leftarrow,\rightarrow,\mathrm{Fire}$ [-1, 1]
Base Env Notes & POMDP Utility RML
Autoencode Push and pop to latent stack, $\text{NC}^{1}$ circuit complexity $O(n)$
CartPole Classic deterministic control task, useful for debugging $O(n)$
CountRecall Increment and query latent counters, $\text{TC}^{0}$ circuit complexity $O(n)$
BattleShip Large configuration space and highly-stochastic transition function $O(n)$
Breakout Disappearing ball, constant RML, and long episodes $O(k)$
MineSweeper Difficult games rules, requires in-context exploration $O(n)$
Navigator Navigation and path planning, requires latent map representation $O(n)$
NoisyPole State estimation under perturbations, surrogate for real-world control $O(n)$
Skittles Obstacle avoidance with flickering obstacles, requires path replanning $O(k)$
Tetris Hard spatial task; human competitions for MDP & POMDP $O(n)$
6 Experiments
Our experiments aim to measure the efficiency of our tasks and analyze memory using our proposed tools. All experiments use the PQN algorithm (Gallici et al., 2024) and five random seeds or five human participants. We describe the experimental setup in Appendices K, L and N and memory models in Appendix J.
Is POPGym Arcade Fast?
We examine the efficiency of our proposed environments. POPGym Arcade uses hardware acceleration for improved throughput, but large pixel observations take significant resources to render. We compare the throughput of our environments to other well-known environments in Fig. 7.
Figure 2: We disentangle the return using our memory analysis tools. We plot the POMDP returns $\in[0,1]$, the Observability Gap, and Memory Bias. We aggregate scores over all environments and difficulty configurations. Whiskers represent the 95% confidence interval over five seeds.
Figure 3: How do policies assign credit? We estimate recall density $\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x},\tau)]$ for the start, middle, and end of a trajectory, aggregating across models and seeds. All density for MDPs should be in $0.66\leq\tau<1.0$. Instead, we see credit diffusely distributed across trajectories across models and tasks.
Figure 4: How do fully-trained agents use memory? We plot pixelwise memory gradients (Eq. 4) for the LRU (top rows) and GRU (bottom rows). In these MDPs, $V_{*}(s_{t})$ is independent of $s_{t-k}\dots s_{t-1}$, yet memory incorrectly smears value credit over uninformative past states, even with a residual connection bypassing the memory model. Smeared value attribution suggests that value estimators may not generalize to new trajectories.
Disentangling Returns
What portion of the return can we attribute to memory models closing the observability gap, and what portion can we attribute to other factors? In Fig. 2, we plot the return, Observability Gap, and Memory Bias aggregated over all tasks and difficulty levels.
How do Agents Use Memory?
We investigate how our trained policies use memory. First, we compute recall densities (Definition 4.3) for each base environment, aggregated over all environment configurations and memory models (Definitions 4.3 and B). Then, we train policies long beyond convergence (Appendix E) and perform qualitative pixel-level gradient visualizations (Eq. 4) for selected configurations (Figs. 4 and C).
Memory Sensitivity Analysis
Are memory-based agents robust to perturbations? We execute toy experiments that introduce Out-of-Distribution (OOD) observations and trajectories into a recurrent policy. First, we collect trajectories following learned policies. Then, we perturb these collected trajectories and analyze how the relative (mean-centered) Q values $A(s,a)=Q(s,a)-|A|^{-1}\sum_{a^{\prime}\in A}Q(s,a^{\prime})$ and policy change. We measure the impact of OOD observations by adding noise to a single frame in each trajectory (Figs. 5 and D). Then, we measure the impact of in-distribution observations but OOD trajectories by shuffling the order of the first ten steps of the trajectory (Fig. 6).
Figure 5: OOD scenarios contaminate the recurrent state and perturb the policy. We add noise $\epsilon$ into one selected observation and examine how this perturbation propagates into current values and actions. The rightmost columns plot the relative $Q$ values (AA), and we color the policy action. We see that not only do the future values change significantly, but so does the policy action.
7 Results
POPGym Arcade is Fast
Our environments are faster than all CPU-based environments we tested (Fig. 7). Our observations are about 10,000 times larger than POPGym, yet we run approximately 100 times faster. We achieve similar throughput to hardware-accelerated Gymnax while generating observations four orders of magnitude larger. One can quickly prototype models or algorithms on our pixel-space tasks.
Figure 6: Even in-distribution observations can contaminate the recurrent state when the trajectory is OOD. We shuffle the beginning of an MDP trajectory and report relative Q values and action. We still see large shifts in value and action. We expect to experience OOD trajectories in offline RL and sim-to-real scenarios.
Cost of Memory
Memory Bias varies between models and can spoil return-only comparisons. For most models, the Gap and Bias have similar scales, demonstrating the weakness of return-only comparisons and the utility of disentangled metrics (Fig. 2). Take the MinGRU and GRU, for example. The relative effect of Bias 0.070.02=0.050.07-0.02=0.05 impacts the return just as strongly as the Observability Gap 0.160.11=0.050.16-0.11=0.05. Purely negative biases across models points to a memory-induced performance penalty, rather than benefit. Although memory models increase parameter count and can approximate a wider class of functions than feedforward networks, such effects are not strong enough to overcome other negative factors. Looking deeper, we note a correlation between the return, Gap, and negative Bias, suggesting a latent factor affecting all three. The following paragraphs hint that optimization difficulty could be one possible culprit, as memory models often learn to construct brittle, overly complex Markov states. However, we cannot prove a causal link between such results and the correlated metrics.
Value Smearing Pathology
Recurrent value functions incorrectly spread credit in a diffuse manner over many or all prior observations, even after converging to optimal or near-optimal returns (Figs. 3 and E). The gradient distributions often appear “smeared” over all prior observations (Figs. 4 and C), even those containing no useful information (e.g., empty board denoting only past actions). We call this phenomena “value smearing”. Value smearing demonstrates that the policies struggle to decorrelate past actions from the forward-looking return, and may signal that the memory model and $Q$ function may have overfit to the trajectory distribution induced by the current policy. We explore this further in the following paragraph.
Recurrent State Contamination
Naturally, we wonder if the reliance on many prior observations means that a single OOD observation can “contaminate” recurrent states and consequently destabilize policies. We find that yes, OOD observations impact the policy far into the future through the recurrent state (Figs. 5 and 14). Even when all observations are in-distribution, OOD trajectories still cause contamination (Fig. 6), although to a lesser extent than OOD observations. These results have far-reaching implications for offline RL, imitation learning, or sim-to-real scenarios, where agents may encounter out-of-distribution scenarios after training.
8 Limitations and Future Work
Understanding Memory Bias
We suggest value smearing as one possible cause for memory-induced performance penalties, but proving causation is difficult and requires future investigation. Future studies should avoid return-only comparisons and utilize the Gap and Bias metrics to measure model capabilities.
MDPs and State Contamination
We highlight some of our findings on MDPs because MDPs provide known ground-truth credit assignment over prior states, while POMDPs do not. We run equivalent experiments on POMDPs and find similar results (Appendix C), but it is difficult to isolate and quantify credit assignment error in POMDPs. Our contamination experiments are proofs of concept, not a rigorous analysis. Non-toy problems have complex observation or transition models that could impact policies differently. We focused on recurrent models primarily due to hardware constraints, but a transformer policy will discard out-of-distribution observations when they exceed the context window, potentially reducing contamination effects.
9 Conclusion
We introduced POPGym Arcade and a methodology to dissect memory in RL, revealing that return-based comparisons can be misleading. We identified a value smearing pathology, where value is incorrectly attributed to irrelevant history. We demonstrated that this credit assignment failure makes memory-endowed policies sensitive to OOD scenarios, corrupting decisions far into the future.
Acknowledgments
This was supported by the University of Macau through SRG2025-00006-FST. Additionally, this work was performed in part at the Super Intelligent Computing Center (SICC) which is supported by IOTSC of the University of Macau.
Reproducibility Statement
We provide memory analysis scripts and full benchmark code in the linked repository in the abstract. We describe memory evaluation tools and POPGym Arcade in main text, while Appendix G gives detailed environment descriptions, Appendix L reports all experiment hyperparameters, and Appendix M outlines our hyperparameter selection methodology. Results are averaged across five random seeds, with additional analyses in Appendices A, B, C, D, E and F. Human baseline procedures are documented in Appendix N, and compute resources are detailed in Appendix O. Together, these materials allow independent researchers to reproduce and extend our findings.
Appendix
Appendix A Environment Throughput
We compare the throughput of our environments to other well-known environments in Fig. 7, using an RTX4090 GPU and Intel I7 13700 CPU. We parallelize CPU environments using synchronous VectorEnvs, and shade the 95% bootstrapped confidence interval.
Figure 7: We plot environment throughput and compare to prior work. We achieve linear scaling until $2^{7}$ environments, after which we saturate GPU cores. POPGym Arcade is faster than Atari, and outperforms POPGym and MinAtar while producing observations orders of magnitude larger.
Appendix B Recall Density Analysis by Model and Task
In this section, we plot the recall density after training, for each model and task. We discretize $\tau$ into sixteen bins and compute the density over five random seeds. No matter the memory model, uninformative prior observations still tend to affect future decisions in MDPs. Even in the best cases, much of the density in MDPs is distributed across old observations (Fig. 10).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 8: Recall density of FART categorized by environment, over five random seeds.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Recall density of GRU categorized by environment, over five random seeds.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 10: Recall density of LRU categorized by environment, over five random seeds.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Recall density of MinGRU categorized by environment, over five random seeds.
Appendix C Additional Pixel-Level Memory Analysis Experiments
We provide further pixel-level analysis of our environment is using Eq. 4. We see that memory models ignore Markov observations, still storing and recalling information in the MDP case.
Figure 12: LRU saliency on the BattleShip task, plotted via Eq. 4 using the L2 norm. The top row represents the MDP and the bottom row represents the equivalent POMDP.
Figure 13: LRU saliency on the Navigator task, plotted via Eq. 4 using the L1 norm. The top row represents the MDP and the bottom row represents the equivalent POMDP.
Appendix D Additional Recurrent State Contamination Experiments
We provide additional experiments for recurrent policy contamination. We use the cartpole example for even longer rollout horizons of 50 and 100 timesteps. We find that old observations still have a significant impact on the relative $Q$ values ($A$) and corresponding actions.
Figure 14: We demonstrate recurrent state contamination over longer horizons, using the LRU model. At 50 and 100 timesteps in the future, corrupted observations can still influence relative $Q$ values ($A$) enough to change the selected action.
Appendix E Pixel Gradient Training Curves
Appendix F Return Analysis by Model
In this section, we provide the unnormalized returns (mean and standard deviation) for the aggregated results in Fig. 2. We provide values for each environment, memory model, and difficulty level.
Table 2: Return for the LRU model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy $0.26\pm 0.00$ $0.27\pm 0.01$
AutoEncodeMedium $0.24\pm 0.01$ $0.24\pm 0.00$
AutoEncodeHard $3\pm 0.02 0.22±0$$2\pm 0.02$
BattleShipEasy $0.85\pm 0.02$ $0.84\pm 0.01$
BattleShipMedium $0.78\pm 0.03$ $0.82\pm 0.01$
BattleShipHard $0.73\pm 0.04$ $0.80\pm 0.01$
BreakoutEasy $0.89\pm 0.01$ $0.75\pm 0.04$
BreakoutMedium $0.83\pm 0.02$ $0.76\pm 0.06$
BreakoutHard $0.84\pm 0.00$ $0.72\pm 0.09$
CartPoleEasy $9\pm 0.00 0.99±0$$9\pm 0.00$
CartPoleMedium $0.98\pm 0.01$ $0.98\pm 0.01$
CartPoleHard $0.97\pm 0.01$ $0.95\pm 0.02$
CountRecallEasy $0.47\pm 0.06$ $0.36\pm 0.01$
CountRecallMedium $0.22\pm 0.03$ $0.23\pm 0.04$
CountRecallHard $0.15\pm 0.02$ $0.23\pm 0.01$
MineSweeperEasy $0.87\pm 0.00$ $0.82\pm 0.01$
MineSweeperMedium $7\pm 0.01 0.66±0$$6\pm 0.01$
MineSweeperHard $0.62\pm 0.01$ $0.61\pm 0.00$
NavigatorEasy $0.91\pm 0.01$ $0.44\pm 0.09$
NavigatorMedium $0.91\pm 0.01$ $0.32\pm 0.03$
NavigatorHard $0.94\pm 0.01$ $0.40\pm 0.06$
NoisyCartPoleEasy $9\pm 0.00 0.99±0$$9\pm 0.00$
NoisyCartPoleMedium $8\pm 0.01 0.99±0$$9\pm 0.00$
NoisyCartPoleHard $0.97\pm 0.01$ $0.97\pm 0.01$
SkittlesEasy $0.83\pm 0.01$ $0.52\pm 0.02$
SkittlesMedium $0.79\pm 0.01$ $0.51\pm 0.00$
SkittlesHard $0.74\pm 0.01$ $0.47\pm 0.00$
TetrisEasy $01\pm 0.00 0.0$$01\pm 0.00$
TetrisMedium $01\pm 0.00 0.0$$01\pm 0.00$
TetrisHard $01\pm 0.00 0.0$$01\pm 0.00$
Table 3: Return for the MinGRU model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy $0.27\pm 0.00$ $0.27\pm 0.00$
AutoEncodeMedium $0.25\pm 0.01$ $0.26\pm 0.00$
AutoEncodeHard $0.25\pm 0.00$ $0.25\pm 0.00$
BattleShipEasy $0.82\pm 0.02$ $0.76\pm 0.01$
BattleShipMedium $0.73\pm 0.02$ $0.75\pm 0.02$
BattleShipHard $0.71\pm 0.03$ $0.73\pm 0.02$
BreakoutEasy $0.88\pm 0.02$ $0.36\pm 0.02$
BreakoutMedium $0.83\pm 0.01$ $0.33\pm 0.04$
BreakoutHard $0.76\pm 0.03$ $0.28\pm 0.02$
CartPoleEasy $0.98\pm 0.01$ $0.87\pm 0.02$
CartPoleMedium $0.93\pm 0.02$ $0.68\pm 0.06$
CartPoleHard $0.88\pm 0.08$ $0.54\pm 0.03$
CountRecallEasy $0.31\pm 0.01$ $0.29\pm 0.03$
CountRecallMedium $0.17\pm 0.00$ $0.21\pm 0.04$
CountRecallHard $0.15\pm 0.01$ $0.13\pm 0.04$
MineSweeperEasy $0.83\pm 0.01$ $0.76\pm 0.01$
MineSweeperMedium $0.64\pm 0.00$ $0.61\pm 0.00$
MineSweeperHard $0.58\pm 0.01$ $0.58\pm 0.01$
NavigatorEasy $0.89\pm 0.02$ $0.33\pm 0.08$
NavigatorMedium $0.74\pm 0.08$ $0.25\pm 0.12$
NavigatorHard $13\pm 0.45 0.0$$09\pm 0.08$
NoisyCartPoleEasy $0.95\pm 0.01$ $0.89\pm 0.01$
NoisyCartPoleMedium $0.92\pm 0.01$ $0.86\pm 0.01$
NoisyCartPoleHard $0.86\pm 0.02$ $0.81\pm 0.01$
SkittlesEasy $0.80\pm 0.00$ $0.39\pm 0.01$
SkittlesMedium $0.76\pm 0.02$ $0.36\pm 0.01$
SkittlesHard $0.73\pm 0.01$ $0.31\pm 0.01$
TetrisEasy $01\pm 0.00 0.0$$01\pm 0.00$
TetrisMedium $\pm 0.01 0.01±0.01$$\pm 0.01$
TetrisHard $1\pm 0.00 0.00±0$$0\pm 0.00$
Table 4: Return for the GRU model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy $0.26\pm 0.00$ $0.26\pm 0.00$
AutoEncodeMedium $0.23\pm 0.01$ $0.24\pm 0.02$
AutoEncodeHard $0.23\pm 0.01$ $0.23\pm 0.02$
BattleShipEasy $0.84\pm 0.01$ $0.84\pm 0.01$
BattleShipMedium $0.76\pm 0.05$ $0.81\pm 0.01$
BattleShipHard $0.71\pm 0.04$ $0.82\pm 0.01$
BreakoutEasy $0.89\pm 0.01$ $0.62\pm 0.10$
BreakoutMedium $0.85\pm 0.01$ $0.60\pm 0.08$
BreakoutHard $0.85\pm 0.02$ $0.44\pm 0.02$
CartPoleEasy $1.00\pm 0.00$ $0.98\pm 0.01$
CartPoleMedium $0.99\pm 0.01$ $0.94\pm 0.02$
CartPoleHard $0.96\pm 0.01$ $0.93\pm 0.04$
CountRecallEasy $0.37\pm 0.04$ $0.36\pm 0.01$
CountRecallMedium $0.19\pm 0.02$ $0.21\pm 0.03$
CountRecallHard $0.16\pm 0.04$ $0.18\pm 0.08$
MineSweeperEasy $0.85\pm 0.01$ $0.81\pm 0.01$
MineSweeperMedium $0.66\pm 0.00$ $0.65\pm 0.01$
MineSweeperHard $0.61\pm 0.01$ $0.61\pm 0.00$
NavigatorEasy $0.90\pm 0.01$ $0.44\pm 0.06$
NavigatorMedium $0.75\pm 0.25$ $0.34\pm 0.09$
NavigatorHard $0.90\pm 0.04$ $0.38\pm 0.06$
NoisyCartPoleEasy $8\pm 0.01 0.99±0$$9\pm 0.00$
NoisyCartPoleMedium $0.97\pm 0.02$ $0.98\pm 0.01$
NoisyCartPoleHard $0.95\pm 0.02$ $0.97\pm 0.02$
SkittlesEasy $0.86\pm 0.01$ $0.50\pm 0.00$
SkittlesMedium $0.83\pm 0.00$ $0.45\pm 0.00$
SkittlesHard $0.75\pm 0.00$ $0.40\pm 0.00$
TetrisEasy $01\pm 0.00 0.0$$01\pm 0.00$
TetrisMedium $01\pm 0.00 0.0$$01\pm 0.00$
TetrisHard $01\pm 0.00 0.0$$01\pm 0.00$
Table 5: Return for the FART model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy $0.25\pm 0.00$ $0.26\pm 0.00$
AutoEncodeMedium $0.23\pm 0.01$ $0.23\pm 0.01$
AutoEncodeHard $0.20\pm 0.03$ $0.21\pm 0.02$
BattleShipEasy $0.66\pm 0.09$ $0.77\pm 0.02$
BattleShipMedium $0.51\pm 0.01$ $0.52\pm 0.02$
BattleShipHard $0.50\pm 0.00$ $0.50\pm 0.00$
BreakoutEasy $0.10\pm 0.04$ $0.13\pm 0.02$
BreakoutMedium $4\pm 0.01 0.11±0$$1\pm 0.05$
BreakoutHard $0.13\pm 0.01$ $0.10\pm 0.05$
CartPoleEasy $0.98\pm 0.01$ $0.97\pm 0.01$
CartPoleMedium $0.85\pm 0.06$ $0.78\pm 0.05$
CartPoleHard $0.62\pm 0.06$ $0.60\pm 0.04$
CountRecallEasy $0.39\pm 0.03$ $0.37\pm 0.02$
CountRecallMedium $0.06\pm 0.00$ $0.10\pm 0.04$
CountRecallHard $06\pm 0.01 0.0$$05\pm 0.00$
MineSweeperEasy $0.76\pm 0.01$ $0.73\pm 0.04$
MineSweeperMedium $0.61\pm 0.02$ $0.59\pm 0.02$
MineSweeperHard $5\pm 0.00 0.55±0$$5\pm 0.01$
NavigatorEasy $0.46\pm 0.14$ $0.36\pm 0.05$
NavigatorMedium $0.27\pm 0.19$ $0.28\pm 0.19$
NavigatorHard $0.45\pm 0.03$ $0.41\pm 0.02$
NoisyCartPoleEasy $0.98\pm 0.01$ $0.97\pm 0.02$
NoisyCartPoleMedium $0.96\pm 0.02$ $0.96\pm 0.01$
NoisyCartPoleHard $0.92\pm 0.02$ $0.89\pm 0.01$
SkittlesEasy $0.85\pm 0.01$ $0.32\pm 0.02$
SkittlesMedium $0.80\pm 0.01$ $0.29\pm 0.01$
SkittlesHard $0.75\pm 0.00$ $0.28\pm 0.01$
TetrisEasy $01\pm 0.00 0.0$$01\pm 0.00$
TetrisMedium $\pm 0.01 0.01±0.01$$\pm 0.01$
TetrisHard $1\pm 0.00 0.00±0$$0\pm 0.00$
Table 6: Return for the MLP model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy $0.26\pm 0.00$ $0.27\pm 0.00$
AutoEncodeMedium $0.25\pm 0.00$ $0.25\pm 0.00$
AutoEncodeHard $0.25\pm 0.00$ $0.25\pm 0.00$
BattleShipEasy $0.85\pm 0.02$ $0.60\pm 0.04$
BattleShipMedium $0.70\pm 0.04$ $0.56\pm 0.01$
BattleShipHard $0.75\pm 0.03$ $0.53\pm 0.00$
BreakoutEasy $0.93\pm 0.01$ $0.49\pm 0.02$
BreakoutMedium $0.89\pm 0.02$ $0.49\pm 0.06$
BreakoutHard $0.84\pm 0.02$ $0.34\pm 0.05$
CartPoleEasy $0.99\pm 0.00$ $0.86\pm 0.02$
CartPoleMedium $0.96\pm 0.01$ $0.47\pm 0.01$
CartPoleHard $0.95\pm 0.01$ $0.31\pm 0.01$
CountRecallEasy $0.40\pm 0.04$ $0.15\pm 0.01$
CountRecallMedium $17\pm 0.02 0.0$$06\pm 0.00$
CountRecallHard $15\pm 0.01 0.0$$08\pm 0.00$
MineSweeperEasy $0.86\pm 0.01$ $0.72\pm 0.01$
MineSweeperMedium $0.65\pm 0.01$ $0.59\pm 0.00$
MineSweeperHard $7\pm 0.00 0.55±0$$5\pm 0.00$
NavigatorEasy $0.88\pm 0.02$ $0.20\pm 0.05$
NavigatorMedium $0.86\pm 0.01$ $0.19\pm 0.02$
NavigatorHard $94\pm 0.01 0.0$$07\pm 0.08$
NoisyCartPoleEasy $0.94\pm 0.01$ $0.82\pm 0.02$
NoisyCartPoleMedium $0.86\pm 0.01$ $0.74\pm 0.01$
NoisyCartPoleHard $0.70\pm 0.01$ $0.66\pm 0.01$
SkittlesEasy $0.84\pm 0.00$ $0.41\pm 0.00$
SkittlesMedium $0.81\pm 0.00$ $0.37\pm 0.00$
SkittlesHard $0.75\pm 0.01$ $0.33\pm 0.00$
TetrisEasy $0.39\pm 0.02$ $0.39\pm 0.02$
TetrisMedium $0.39\pm 0.02$ $0.39\pm 0.02$
TetrisHard $2\pm 0.01 0.22±0$$2\pm 0.01$
Appendix G Environment Descriptions
We describe each of our implemented MDPs and rules. Then, we explain how and why we make each MDP into a POMDP.
G.1 Battleship
MDP
The agent’s goal is to sink all the pre-placed ships to win the game. The agent controls a cursor that represents its current position on the grid. Each turn, the agent can either move the cursor one space or choose to ”FIRE” the current grid cell. Each cell can be in one of three states: COVERED, HIT, or MISS. Initially, all cells are COVERED, but each changes to HIT or MISS upon firing. The agent receives a positive reward for hitting a ship, a neutral reward (0) for the first miss on a COVERED cell, and a negative reward for repeatedly firing HIT or MISS cells.
POMDP
We make this task partially observable by only showing the HIT or MISS markers for the tile the cursor currently occupies. All other tiles appear COVERED, regardless of whether they have been fired upon. This does not affect the reward function, only the observation function.
POMDP Required Capabilities
The agent must remember the tile label (HIT/MISS) of previously fired upon tiles, or risk running out of moves. This capability is critical for avoiding redundant actions, which incur negative rewards, and performing an efficient search of locations and ships. This will test spatial memory and integrate a sequence of localized observations into a coherent ability.
MDP Recovery from POMDP
By retaining all prior observations, the agent can view all the same HIT/MISS markers as in the MDP.
G.2 Count Recall
MDP
Every turn, the agent gets a value card and a query card. All previous value cards are displayed to the agent. The agent’s task is to figure out how many times the query card has shown up so far. To do this, the agent counts how many times it has received the matching value card and uses that as the answer. If the agent guesses the correct count, it receives a positive reward. If the guess is incorrect, the agent receives no reward (0). This setup encourages the agent to accurately track and recall the frequency of specific cards over time.
POMDP
All previous value cards are hidden, the agent may only observe the current value and query card.
POMDP Required Capabilities
The agent must learn to implement a latent counter for each card type to track the number of times a card has appeared. At each timestep, the agent must utilize these counters to answer the query.
MDP Recovery from POMDP
By retaining all prior observations, the agent can reconstruct the MDP view from the suits at the top of the screen.
G.3 Mine Sweeper
MDP
The agent’s goal is to hit all the non-mine cells to win the game, inspired by the classic computer game Minesweeper. Similar to Battleship environments, the agent controls a cursor to navigate the grid and can either move to an adjacent cell or sweep the current cell each turn. Sweeping a safe cell earns the agent a positive reward and reveals the number of adjacent mines, while hitting a mine results in negative reward and ends the game. If the agent tries to sweep a cell that has already been revealed, it receives a small negative reward. This reward structure encourages the agent to explore efficiently, avoid mines, and minimize redundant actions to maximize its cumulative reward.
POMDP
In the POMDP variant, the agent can only observe the tile at the cursor position. All other tiles appear covered.
POMDP Required Capabilities
The agent must remember the number of adjacent mines for previously swept tiles, while simultaneously learning the fairly complex rules of the game. In particular, the agent must predict the location of mines from a partial view of adjacency information, which itself must be memorized.
MDP Recovery from POMDP
By retaining all prior observations, the agent can view all the same tile values as in the MDP.
G.4 Autoencode
MDP
This game is similar to Simon but played in reverse. The agent’s task is to recall and replay a sequence of cards in the opposite order they were shown. During the WATCH phase, the agent observes a randomly generated sequence of cards. This sequence is displayed on-screen. In the PLAY phase, the agent must reproduce the sequence in reverse order. The agent receives a positive reward for each correct card played and no reward (0) for incorrect choices.
POMDP
This variant does not display the series of cards to the screen. At each timestep during the WATCH phase, the agent receives only the corresponding card.
POMDP Required Capabilities
By playing the cards in reverse order, the agent must learn to push and pop from a stack in latent space. Sequence permutation (pushing and popping) is well-known to be $\text{NC}^{1}$ complexity and not solvable by transformers in constant depth (Merrill et al., 2024). Only nonlinear RNNs can solve such a task in constant depth.
MDP Recovery from POMDP
By retaining all prior observations, the agent can reconstruct the MDP view using the suits at the top of the screen.
G.5 Navigator
MDP
The agent’s goal is to navigate and open the treasure on the board while avoiding the TNT. The agent navigates using a cursor that marks its current position. On each turn, it can either shift the cursor to an adjacent cell or decide to open the cell it is currently in. Every move the agent makes results in a small negative reward. If the agent lands on a TNT cell, it receives a significant negative reward, and the game ends immediately. This environment challenges the agent to discover the most efficient path to the treasure, aiming to maximize its cumulative reward by completing the game as quickly as possible.
POMDP
The agent only sees the position of the treasure and TNT blocks at the initial timestep. Afterwards, the agent can only see its current position.
POMDP Required Capabilities
The agent must remember the position of the treasure and TNT blocks over time. It must perform path planning within the learned memory state to reach the goal quickly.
MDP Recovery from POMDP
By retaining the first observation, the agent has access to the goal location. Retaining all consecutive actions enables reconstruction of the agent’s current location.
G.6 Skittles
MDP
This game is inspired by many Atari games that focus on moving the main character to avoid touching enemies. The agent is a white square at the bottom of the screen and must dodge colored falling blocks. Upon touching a colored block, the game terminates. The agent receives a positive reward for surviving and a negative reward for touching a colored block. We spawn the colored blocks in such a way that the agent always has a feasible path for survival.
POMDP
In the POMDP variant, each colored block has only a 50% probabilty of being rendered at each timestep, causing them to flicker in and our of the agent’s view, the agent cannot rely on the current observation to navigate safely.
POMDP Required Capabilities
The agent must use its memory to develop a form of object permanence, tracking the downward trajectories of blocks even when they are temporarily invisible. This requires the agent to maintain an internal belief about all colored blocks, integrating history observations to infer the complete observation of the envvironment.
MDP Recovery from POMDP
The blocks have a 50% chance of rendering. Given that there are 11 tiles between the block spawn point and the agent, there is a $1-0.5^{11}=0.9995$ chance that each block will be rendered at least once. By retaining all frames, the agent can reconstruct the position of all blocks (MDP state).
G.7 Breakout
MDP
This game is based on the classic Atari Breakout game. The agent controls a paddle that they may use to deflect a moving ball. When the ball touches colored blocks, the blocks deflect the ball and then disappear. The agent receives positive reward when the ball collides with colored blocks, and a negative reward and termination condition when the ball passes the paddle and leaves the bottom of the play area. To make this game fully observable, we also provide a ball “tail” that determines the velocity direction and magnitude of the ball. Upon clearing all the blocks, the game terminates.
POMDP
In the POMDP variant, we remove the tail. Furthermore, the ball is only visible when moving upward. It becomes invisible when moving downward. We make sure to start the game with the ball moving upward so the agent knows the initial position and velocity of the ball.
POMDP Required Capabilities
Unlike other tasks, Breakout enables episodes that are thousands of timesteps long, allowing us to understand how memory adapts over very long sequences. This POMDP primarily tests short-term memory, as the agent need only (1) integrate position to predict velocity and (2) remember where the ball was a few timesteps ago to predict the future path of the ball. But it must learn a representation that can do so reliably over long durations.
MDP Recovery from POMDP
By retaining all observations up to and including collision of the ball with the block, the agent can predict the deterministic angle and velocity at which the ball will reflect. This information is sufficient to predict the position and velocity of the ball, recovering the MDP.
G.8 Tetris
MDP
This is the well-known Russian puzzle game. Colored blocks of various shapes are dropped from the top of the screen, and the player controls their descent (position and rotation). Upon reaching the bottom of the screen or touching another block, the controlled block is frozen in-place and a new agent-controlled block is spawned. The blocks slowly pile up, and the agent receives a negative reward and termination upon the blocks reaching the top of the screen. Completely filling a row with blocks “clears” the row, deleting the tiles in the row and providing a positive reward. The game also terminates after clearing a predetermined number of rows.
POMDP
In the POMDP variant, once a block is frozen in place, it becomes invisible. Only the currently controlled block and block-clear animation are visible to the agent. This is known as “Invisibile Tetris” and has a significant human following and some associated human competitions as well.
POMDP Required Capabilities
Invisible Tetris is arguably the most difficult task we propose. Standard Tetris is already a fairly hard task, as as far as we know, unsolved by RL. Even Tetris Grand Masters struggle with Invisible Tetris. The arrangement of tiles is complex and constantly shifting. It must be memorized perfectly, and a small error in state quickly compounds as the hidden structure undergoes mutation. This POMDP requires very strong state tracking capabilities and the ability to learn very difficult games.
MDP Recovery from POMDP
Each block is visible during the timestep it is locked in place. Considering all observations therefore provides all block positions and recovers the MDP.
G.9 CartPole
MDP
The CartPole environment is a classic control problem framed as a Markov Decision Process (MDP), where the objective is to balance a pole on a moving cart.
POMDP
In the POMDP version, after the initial timestep, the cart and pole are hidden from the agent while the arrows representing velocity remain. A horizontal arrow representing the cart’s velocity and a curved arrow representing the pole’s angular velocity. The size of each arrow is proportional to the magnitude of the corresponding velocity. This creates a pixel-form of the position-masked CartPole problem.
POMDP Required Capabilities
To succeed, the agent must integrate velocity signals from all prior timesteps to infer the hidden positional information, directly testing the model’s capacity to maintain an internal state from a stream of partial information.
MDP Recovery from POMDP
By considering the initial position (rendered) and integrating over all velocities, the agent can predict cart and pole position information.
G.10 NoisyPole
MDP
This task is Cartpole (Env. 1) affected by Gaussian noise on position and speed of the cart as well as angle and angular velocity of the pole. The result of noise is still reflected in the arrow magnitudes.
POMDP
After the initial timestep, the cart and pole are hidden from the agent, providing only the velocity arrows.
POMDP Required Capabilities
This is a harder version of the CartPole POMDP. Morad et al. (2023a) demonstrates that certain memory-free policies can do well in POMDP CartPole, perhaps by learning a time-varying policy. Adding noise results in a better test of memory, as the agent must remember and react to the noise. A simple time-varying policy is insufficient to solve this task.
MDP Recovery from POMDP
By considering the initial position (rendered) and integrating over all velocities, the agent can predict cart and pole position information. Noise information is reflected in the velocity observations, allowing integration to recover the full state.
Appendix H Hardware-Accelerated RL
Copying observations from CPU to GPU during rollouts is a major efficiency bottleneck in the training process (Lu et al., 2022). Using JAX, we can implement environments on accelerator hardware, avoiding copy overhead and training policies up to one thousand times faster (Lange, 2022; Koyamada et al., 2023; Matthews et al., 2024b). We highlight Matthews et al. (2024a); Pignatelli et al. (2024); Lu et al. (2024) which offer hardware-accelerated POMDPs. With improved environment throughput, we can consider new training paradigms like Podracer (Hessel et al., 2021; Toledo, 2024). In this work, we focus on the PQN algorithm (Gallici et al., 2024), a simplified podracer version of Q learning. Unlike DQN (Mnih et al., 2015), PQN does not use target networks or replay buffers, and opts for an on-policy TD($\lambda$) objective.
Given the poor sample efficiency of RL, we focus on maintaining high environment throughput. First and foremost, we ensure that all environments are implemented in JAX (Bradbury et al., 2018) in a vectorizable and compilable form to leverage hardware acceleration. Beyond this, we find a number of small tricks can improve performance. Pre-caching sprites and their pixel positions upon reset provides noticeable performance gains, and replacing calls to jax.lax.cond with jax.lax.switch provides increased throughput.
H.1 Vectorized State Transitions and Rendering
Our rendering system is built on a canvas framework. We draw to the canvas using primitives, such as letters, numbers, shapes, and more. All rendering functions are pure (without side effects), and may be easily vectorized or compiled. We build our rendering framework on top of jax.numpy and does not use any external dependencies. With these tools, anyone can easily render their own custom grid-based or card-based environments for further research and exploration.
Seven of ten of our proposed environments come with fairly intricate rules, requiring multiple iterations of pattern rendering on the canvas. These complex tasks can be broadly split into two rendering categories: grid-based environments, like Battleship, and card-based environments, like CountRecall. Initially, we re-rendered at each step, but this was slow, even with JIT compilation. We found that removing dynamic computations improved performance, which we detail below.
Grid-Based Environments
In grid-based games, we divide the entire canvas into a bunch of small square cells, where each cell displays either the same or a unique pattern to reflect the environmental info contained in the current state. We precompute the coordinates of every cell that needs to be drawn on the canvas and stash them in two matrices – one for the x-axis positions and one for the y-axis. During rendering, we leverage jax.vmap to handle the process in parallel along the row dimension.
Card-Based Environments
We precompute all possible card templates (value, query, and historical cards) during initialization period, avoiding repetitive redrawing of static elements like suits or card positions. This turns dynamic rendering into a fast lookup-and-merge process. For example, when rendering value or query cards, we retrieve pre-drawn templates and apply them to the canvas using masks, skipping pixel-by-pixel logic during runtime. For the history display, we vectorize the drawing of historical cards using jax.vmap, generating all variations upfront and later selecting only the visible ones with fast array indexing. By JIT-compiling the render function, we lock these optimizations into a highly efficient, hardware-accelerated pipeline. The benefits are reduced per-frame computation, minimal branching, and GPU-friendly operations—all critical for real-time rendering. Even when rendering complex elements like the history grid, we avoid loops and instead use clever indexing with jnp.argmax and masks to overlay the latest valid symbols.
Appendix I Related Benchmarks
Below, we list additional relevant benchmarks and their characteristics. A tilde in the MDP Twins columns means that MDP and POMDP twins do not share observation spaces, making causal studies difficult. Our benchmark is the only one with POMDP/MDP twins and GPU acceleration.
Benchmark POMDP/MDP Twins GPU
MuJoCo/MJX (Todorov et al., 2012)
Atari/ALE (Bellemare et al., 2013)
DMLab (Beattie et al., 2016)
MiniGrid/Navix (Chevalier-Boisvert et al., 2018), (Pignatelli et al., 2024) $\sim$
Memory Task Suite (Fortunato et al., 2019)
MinAtar (Young & Tian, 2019)
EnvPool Weng et al. (2022)
Gymnax (Lange, 2022)
POMDP Baselines (Ni et al., 2022) $\sim$
MemoryGym (Pleines et al., 2022)
POPGym/POPJaxRL (Morad et al., 2023a), (Lu et al., 2024) $\sim$
PGX (Koyamada et al., 2023)
MDP Playground (Rajan et al., 2023)
Jumanji (Bonnet et al., 2024)
Kinetix (Matthews et al., 2024b)
Craftax (Matthews et al., 2024a)
POBAX (Tao et al., 2025) $\checkmark$
POPGym Arcade (ours)
Appendix J Recurrent Models
In this appendix, we provide detailed descriptions of the recurrent models used in our experiments. We categorize these models into two families: classical recurrences, such as LSTMs and GRUs, which process sequences step-by-step, and associative recurrences, which leverage parallelizable operators for significantly faster computation.
J.1 Classical Recurrences
Classical recurrent models, such as the Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU), process sequential information in a strictly ordered manner. The computation of the hidden state at any step $t$, denoted as $h_{t}=f(h_{t-1},x_{t})$, is fundamentally dependent on the completion of the previous step’s computation, $h_{t-1}$. This creates a sequential chain of operations that cannot be parallelized.
LSTM
The LSTM network (Hochreiter & Schmidhuber, 1997) introduces a dedicated cell state $c_{t}$ to carry information over long sequences, separate from the hidden state $h_{t}$. Information flow is regulated by three gates: an input gate $i_{t}$, a forget gate $f_{t}$, and an output gate $o_{t}$. At each timestep $t$, the LSTM updates its states as follows:
$f_{t}$ $=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$ (9)
$i_{t}$ $=\sigma(W_{i}[h_{t-1},x_{t}]+b_{i})$ (10)
$o_{t}$ $=\sigma(W_{o}[h_{t-1},x_{t}]+b_{o})$ (11)
$\tilde{c}_{t}$ $=\tanh(W_{c}[h_{t-1},x_{t}]+b_{c})$ (12)
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}_{t}$ (13)
$h_{t}$ $=o_{t}\odot\tanh(c_{t}),$ (14)
where $\sigma$ is the sigmoid function, $\odot$ denotes element-wise multiplication, and $h_{t-1},x_{t}]$ is the concatenation of the previous hidden state and the current input.
GRU
The Gated Recurrent Unit (GRU) (Chung et al., 2014) simplifies the LSTM architecture by merging the cell state and hidden state into a single hidden state vector $h_{t}$. It uses two gates: a reset gate $r_{t}$ and an update gate $z_{t}$. The reset gate determines how to combine the new input with the previous hidden state, while the update gate decides how much of the previous hidden state to retain. The update equations are:
$z_{t}$ $=\sigma(W_{z}[h_{t-1},x_{t}]+b_{z})$ (15)
$r_{t}$ $=\sigma(W_{r}[h_{t-1},x_{t}]+b_{r})$ (16)
$\tilde{h}_{t}$ $=\tanh(W_{h}[r_{t}\odot h_{t-1},x_{t}]+b_{h})$ (17)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}.$ (18)
The GRU’s simpler structure makes it computationally more efficient than the LSTM while often achieving comparable performance.
J.2 Associative Recurrences
Associative recurrent models, also known as linear recurrent models or memoroids (Morad et al., 2024), update the recurrent state with a binary operator $\bullet$ that obeys the associative property
$h_{3}=(x_{1}\bullet x_{2})\bullet x_{3}=x_{1}\bullet(x_{2}\bullet x_{3}).$ (19)
Given an associative recurrent update, we can leverage the associative property to rearrange the order of operations. For example, given four inputs $x_{1},x_{2},x_{3},x_{4}$, we can compute either
$h_{4}=(((x_{1}\bullet x_{2})\bullet x_{3})\bullet x_{4})$ $h_{4}=(x_{1}\bullet x_{2})\bullet(x_{3}\bullet x_{4}).$ (20)
The former case corresponds to standard recurrent networks, relying on the prior recurrent state to compute the current state, resulting in linear time complexity. In the latter case, we can compute $x_{1}\bullet x_{2}$ and $x_{3}\bullet$ $x_{4}$ in parallel, achieving logarithmic parallel time complexity (Hinze, 2004). Blelloch (1990) provide an associative scan implementation with linear space complexity. Associative recurrences are orders of magnitude faster than classical RNNs in practice, while using much less memory than transformers, making them useful in sample inefficient tasks like RL (Lu et al., 2024).
Our experiments rely on three distinct associative recurrent models: the Fast Autoregressive Transformer (FART) (Katharopoulos et al., 2020), a form of State-Space Model (Gu et al., 2022) called the Linear Recurrent Unit (LRU)(Orvieto et al., 2023), and a GRU (Chung et al., 2014) variant of a Minimal Recurrent Network called the MinGRU (Feng et al., 2024). We provide formal descriptions of each model below.
Linear Transformers
Standard transformers have quadratic space complexity from the outer product of keys and queries. The Fast Autoregressive Transformer (FART) (Katharopoulos et al., 2020) replaces softmax attention with a kernelized attention mechanism to achieve linear space complexity and logarithmic time complexity via associative recurrent updates
$h_{t}=h_{t-1}+\phi(W_{k}x_{t})\phi(W_{v}x_{t})^{\top}$ $\hat{s}_{t}=\mathrm{MLP}\left(x+\frac{\phi(W_{q}x_{t})^{\top}h_{t}}{\phi(W_{q}x_{t})\cdot\sum_{i=0}^{t}\phi(W_{k}x_{i})}\right).$ (21)
Here, $\phi(x)=1+\mathrm{ELU}(x)$ represents a kernel-space projection and the recurrent state $h$ represents the attention matrix. $W_{k},W_{v},W_{q}$ represent the key, query, and value projection parameters. To compute $\hat{s}_{t}$, we multiply the recurrent attention matrix by the query vector and normalize by a scalar.
State-Space Models
State-Space Models (SSMs) (Gu et al., 2022) model an associative recurrence via
$h_{t}=\overline{W}_{A}h_{t-1}+\overline{W}_{B}x_{t}$ $\hat{s}_{t}=\overline{W}_{C}h_{t}+\overline{W}_{D}x_{t},$ (22)
where $\overline{W}_{A},\overline{W}_{B},\overline{W}_{C},\overline{W}_{D}$ are discretizations of carefully initialized trainable parameters $W_{A},W_{B},W_{C},W_{D}$. In practice, we initialize $W_{A},W_{B},W_{C},W_{D}$ deterministically. Deterministic initialization applied to many consecutive SSM layers can result in instabilities, and so Orvieto et al. (2023) proposes a meticulously derived random initialization and new parameterization of SSM parameters. They call their method the Linear Recurrent Unit (LRU).
Minimal Recurrent Networks
Feng et al. (2024) revisit the popular Gated Recurrent Unit (GRU) (Chung et al., 2014) and Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) RNNs. They simplify the GRU and LSTM recurrent updates, proposing the MinGRU and MinLSTM with efficient associative recurrent updates. The authors write the MinGRU as
$h_{t}=(1-\sigma(W_{1}x_{t}+b_{1}))\odot h_{t-1}+\sigma(W_{1}x_{t}+b_{1})\odot\tanh(W_{2}x_{t}+b_{2})$ $\hat{s}_{t}=h_{t},$ (23)
with trainable parameters $W_{1},b_{1},W_{2},b_{2}$, sigmoid function $\sigma$, and elementwise product $\odot$. The authors find that the MinGRU outperforms multiple SSM variants across offline reinforcement learning tasks, using an offline decision transformer (Chen et al., 2021) framework.
Appendix K Network Architecture
PQN
Our Q-network combines spatial feature extraction with decision policy learning through a JAX/Equinox implementation. The network processes batches of $128\times 128$ RGB images through four convolutional blocks, followed by three dense layers with intermediate normalization (Table 7).
Table 7: Q network architecture
Layer Parameters Activation
Conv2D-1 Channels: $3,64$, Kernel: $5\times 5$, Stride: 2 LeakyReLU
MaxPool2D-1 Pool: $2\times 2$, Stride: 2
Conv2D-2 Channels: $64,128$, Kernel: $3\times 3$, Stride: 2 LeakyReLU
MaxPool2D-2 Pool: $2\times 2$, Stride: 2
Conv2D-3 Channels: $128,256$, Kernel: $3\times 3$, Stride: 2 LeakyReLU
MaxPool2D-3 Pool: $2\times 2$, Stride: 2
Conv2D-4 Channels: $256,512$, Kernel: $1\times 1$, Stride: 1 LeakyReLU
Dense-1 Features: $512,256$ LeakyReLU
LayerNorm Shape: 256
Dense-2 Features: $256,256$ LeakyReLU
LayerNorm Shape: 256
Dense-3 Features: $256,5$
PQN $\times$ RNN
Our recurrent Q-network extends the base architecture with explicit memory handling through a hybrid CNN-RNN-MLP design. The RNN combines a 512-channel input tensor x with a 5-dimensional one-hot encoded last action vector from POPGym Arcade, processes these through its 512-unit hidden state, and generates 256-dimensional output features (Table 8).
Table 8: Recurrent Q network architecture
Layer Parameters Activation
Conv2D-1 Channels: $3,64$, Kernel: $5\times 5$, Stride: 2 LeakyReLU
MaxPool2D-1 Pool: $2\times 2$, Stride: 2
Conv2D-2 Channels: $64,128$, Kernel: $3\times 3$, Stride: 2 LeakyReLU
MaxPool2D-2 Pool: $2\times 2$, Stride: 2
Conv2D-3 Channels: $128,256$, Kernel: $3\times 3$, Stride: 2 LeakyReLU
MaxPool2D-3 Pool: $2\times 2$, Stride: 2
Conv2D-4 Channels: $256,512$, Kernel: $1\times 1$, Stride: 1 LeakyReLU
RNN Cell Input, Hidden, Output: $517,512,256$, Num Layer: 2
Dense Features: $256,5$
Appendix L Experiment Hyperparameters
We used one set of hyperparameters for all our experiments. Please see the following section for the hyperparameter selection methodology.
Table 9: PQN hyperparameters used in all of our experiments. See Gallici et al. (2024) for a detailed description of hyperparameters.
Parameter Value
TOTAL_TIMESTEPS 10e6, 20e6
TOTAL_TIMESTEPS_DECAY 1e6, 2e6
NUM_ENVS 16
NUM_STEPS 128
EPS_START 1.0
EPS_FINISH 0.05
EPS_DECAY 0.25
NUM_MINIBATCHES 16
NUM_EPOCHS 4
NORM_INPUT False
NORM_TYPE layer norm
LR 0.00005
MAX_GRAD_NORM 0.5
LR_LINEAR_DECAY True
REW_SCALE 1.0
GAMMA 0.99
LAMBDA 0.95
Appendix M Hyperparameter Selection Methodology
To select hyperparameters, we performed a manual sweep over four model architectures (MLP, MinGRU, LRU, and FART) across six tasks (CartPole, Navigator, BattleShip, MineSweeper, CountRecall, and AutoEncode), and all three difficulty levels. We evaluated the training timesteps, learning rate and schedule, exploration parameters, batch size, number of minibatches per epoch, gradient clipping magnitude. These refer to TOTAL_TIMESTEPS, LR, LR_LINEAR_DECAY, TOTAL_TIMESTEPS_DECAY, EPS_START, EPS_FINISH, EPS_DECAY, NUM_STEPS, NUM_ENVS, NUM_MINIBATCHES, MAX_GRAD_NORM respectively.
Unlike the original PQN paper which annealed learning epsilon to near-zero over the entire training duration, we found it beneficial to decay more quickly to a slightly higher final epsilon value. When experimenting with epsilon decay to $0.01$ and $0.05$ over the episode, we noticed that most learning tended to happen near the ends of training. By quickly annealing to $0.05$, the policies learned much more quickly. We found annealing epsilon to $0.01$ produced suboptimal results.
The batch size is a function of the number of workers and the number of steps taken at each epoch. We found decreasing the number of steps below 128 hurt performance, and increasing it beyond 128 yielded similar results but reduced sample efficiency. With a number of steps at 128, we found that doubling the batch size via double the number of workers, did not produce a noticeable improvement. Performance increased as we increased the number of minibatches to 16, beyond which we did not see meaningful improvements.
We found that changing lambda from the PQN recommended 0.65 to our selected 0.95 resulted in the best returns. We evaluated learning rates including $0.0005,0.0001,0.00005,0.00001$ finally settling on $0.00005$. We found linear LR decay outperformed no decay, and that decaying over the first tenth of an episode produced better results than decaying over the full training duration. We tested gradient clipping values of 1.0 and 0.5, selecting the smaller value for training stability despite a minor efficiency reduction.
Appendix N Human Baselines
We added human baseline results in our study. Each human was provided the docstring associated with each game, then played the game using the arrow keys and spacebar. The scores were recorded and uploaded for our analysis.
These baselines are computed from five participants, matched against the five seeds reported in the Fig. 2. We report the per-task breakdown of normalized returns in $[0,1]$ at the end of this response.
In general, humans tend to perform better on card games, perform similarly on board games, and perform worse on control tasks. Overall, the MLP outperforms humans on MDPs. Humans make up this gap when introducing partial observability, scoring almost the same as the MLP.
Table 10: Return for the human baselines, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy $0.79\pm 0.39$ $0.28\pm 0.09$
AutoEncodeMedium $0.82\pm 0.17$ $0.27\pm 0.08$
AutoEncodeHard $0.58\pm 0.12$ $0.28\pm 0.04$
BattleShipEasy $0.96\pm 0.09$ $0.71\pm 0.08$
BattleShipMedium $0.92\pm 0.08$ $0.73\pm 0.29$
BattleShipHard $0.94\pm 0.13$ $0.66\pm 0.22$
CartPoleEasy $0.22\pm 0.11$ $0.18\pm 0.10$
CartPoleMedium $16\pm 0.10 0.0$$07\pm 0.05$
CartPoleHard $12\pm 0.12 0.0$$07\pm 0.05$
CountRecallEasy $0.78\pm 0.10$ $0.35\pm 0.20$
CountRecallMedium $0.68\pm 0.17$ $0.15\pm 0.08$
CountRecallHard $0.57\pm 0.11$ $0.14\pm 0.09$
MineSweeperEasy $0.78\pm 0.16$ $0.79\pm 0.13$
MineSweeperMedium $0.41\pm 0.21$ $0.30\pm 0.14$
MineSweeperHard $0.26\pm 0.28$ $0.17\pm 0.17$
NavigatorEasy $0.97\pm 0.02$ $0.97\pm 0.02$
NavigatorMedium $0.97\pm 0.01$ $0.97\pm 0.01$
NavigatorHard $0.97\pm 0.01$ $0.68\pm 0.67$
NoisyCartPoleEasy $0.09\pm 0.05$ $0.16\pm 0.12$
NoisyCartPoleMedium $0.19\pm 0.13$ $0.17\pm 0.08$
NoisyCartPoleHard $0.18\pm 0.09$ $0.14\pm 0.12$
BreakoutEasy $55\pm 0.45 0.0$$08\pm 0.03$
BreakoutMedium $81\pm 0.39 0.0$$06\pm 0.02$
BreakoutHard $72\pm 0.39 0.0$$06\pm 0.05$
SkittlesEasy $0.80\pm 0.33$ $0.54\pm 0.33$
SkittlesMedium $0.76\pm 0.32$ $0.59\pm 0.39$
SkittlesHard $0.69\pm 0.30$ $0.42\pm 0.24$
TetrisEasy $21\pm 0.14 0.0$$02\pm 0.03$
TetrisMedium $18\pm 0.12 0.0$$03\pm 0.03$
TetrisHard $3\pm 0.04 0.00±0$$0\pm 0.00$
Appendix O Compute Resources
We used approximately 231 days equivalent training time on an RTX4090 GPU to produce a little over 2,000 total runs. We used an unknown amount of additional resources to tune hyperparameters and rerun experiments after fixing bugs.
Appendix P LLM Usage
We used LLMs as a writing aid for portions of the paper. We used it to improve grammar and clarity, as well as to restructure the order in which we present our ideas.